# AthletePose3D: A Benchmark Dataset for 3D Human Pose Estimation and Kinematic Validation in Athletic Movements

### v1

@@ 2 Related work
Large-scale 3D human pose datasets have been fundamental for training and evaluating monocular 3D pose estimation models. Table 1 provides a detailed summary of existing 3D human pose datasets. Significant efforts have been made to develop large-scale datasets, with Human3.6M [21] and MPI-INF-3DHP [45] being among the most well-known benchmarks for 3D pose estimation. Broadly, 3D human pose datasets are collected using two mainstream methods: marker-based motion capture, which employs optical tracking systems and infrared cameras (e.g., Human3.6M [21], 3DWP [70], HumanEva [58], TotalCapture [68]), and markerless motion capture, which relies on industrial-grade cameras and computer vision algorithms (e.g., MPI-INF-3DHP [45], CMU Panoptic [25], SportsPose [20], FS-Jump3D [63], WorldPose [24]).
@@
For 3D human posture in sports, SportsPose [20] and ASPset-510 [49] are currently the largest datasets designed to facilitate 3D human posture estimation. Other datasets, such as 3DSP [80] and WorldPose [24], provide posture data from professional soccer matches for posture analysis and pose tracking, respectively, while FS-Jump3D [63] focuses on figure skating postures to support temporal action segmentation.
However, existing datasets are often not specifically designed for 3D pose estimation or lack high-speed, high-acceleration movements that are essential in competitive sports. To address these limitations, we introduce AthletePose3D, a comprehensive dataset encompassing 12 types of sports actions performed by athletes ranging from amateurs (athletes without remuneration) to professionals. See Section 4 for more details.
@@ 2.2 Monocular 2D & 3D human pose estimations
Monocular 3D pose estimation methods can be broadly classified into direct 3D pose estimation [8, 22, 54] and 2D-to-3D lifting [44, 79, 60], where the latter involves reconstructing 3D poses from 2D keypoint coordinates. Early approaches to 3D pose estimation were based on Hidden Markov Models [27, 67] and graphical models [26, 59]. With the rise of deep learning, Fully Connected Networks [11, 43] were introduced, followed by temporal-based [7, 50] and graph-based [2, 9, 71] convolutional networks, which improved the modeling of sequential dependencies and inter-keypoint relationships.
More recently, transformer-based methods have become the dominant approach, achieving state-of-the-art performance. PoseFormer [85] pioneered this direction, leading to further advancements with models such as [84, 34, 82, 64, 57, 56, 13, 33, 64, 52]. The latest models, KTPFormer [51] integrates kinematics and trajectory priors to improve motion representation, MotionAGFormer [44] introduces a hybrid approach, combining self-attention with graph convolution to better capture both spatial and temporal dependencies, and TCPFormer [37] further enhances temporal modeling by leveraging pose proxies to refine sequential feature learning. Additionally, self-supervised learning has been extensively explored in 3D pose estimation [30, 72, 14, 48, 78, 86].
Although many monocular pose estimation models have been proposed, they are often benchmarked on datasets featuring everyday movements. High-speed sports movements, however, present unique challenges, and the performance of these models on such movements remains unclear. Therefore, in this study, we select the best-performing 2D & 3D models from corresponding benchmark datasets and fine-tune them on the AthletePose3D dataset to evaluate their performance.
@@ 2.3 Validate kinematic from pose estimation
Although models like OpenPose [4], Mediapipe [41], and HRNet [60] are frequently used in sports science and kinematic validation studies, they do not represent the state-of-the-art in 2D pose estimation. Moreover, 3D pose estimation models have received limited attention in validation studies. Therefore, this study evaluates both state-of-the-art 2D and 3D pose estimation models for sports motion kinematic waveform using the newly introduced AthletePose3D dataset.
@@ 3.2 Benchmarking monocular 2D & 3D pose estimation models
To benchmark both monocular 2D and 3D pose estimation models, we considered the best-performing models on representative datasets, taking into account their architectures. These models were fine-tuned and evaluated using the AthletePose3D dataset, which was split into training, validation, and test sets with a 60/20/20 ratio. The training set was used for model training, and the validation set results were reported..

### v2

@@ 2 Related work
Large-scale 3D human pose datasets have been fundamental for training and evaluating monocular 3D pose estimation models. Table 1 provides a detailed summary of existing 3D human pose datasets. Significant efforts have been made to develop large-scale datasets, with Human3.6M [21] and MPI-INF-3DHP [45] being among the most well-known benchmarks for monocular 3D pose estimation. Broadly, 3D human pose datasets are collected using two mainstream methods: marker-based motion capture, which employs optical tracking systems and infrared cameras (e.g., Human3.6M [21], 3DWP [70], HumanEva [58], TotalCapture [68]), and markerless motion capture, which relies on industrial-grade cameras and computer vision algorithms (e.g., MPI-INF-3DHP [45], CMU Panoptic [25], SportsPose [20], FS-Jump3D [63], WorldPose [24]).
@@
For 3D human posture in sports, SportsPose [20] and ASPset-510 [49] are currently the largest datasets designed to facilitate monocular 3D human posture estimation. Other datasets, such as 3DSP [80] and WorldPose [24], provide posture data from professional soccer matches for posture analysis and pose tracking, respectively, while FS-Jump3D [63] focuses on enhancing figure skating 3D pose estimation.
However, existing datasets are often not specifically designed for monocular 3D pose estimation or lack high-speed, high-acceleration movements that are essential in competitive sports. To address these limitations, we introduce AthletePose3D, a comprehensive dataset encompassing 12 types of sports actions performed by athletes ranging from amateurs (athletes without remuneration) to professionals. See Section 4 for more details.
@@ 2.2 Monocular 2D & 3D human pose estimations
Monocular 3D pose estimation methods can be broadly classified into direct 3D pose estimation [8, 22, 54] and 2D-to-3D lifting [44, 79, 60], where the latter involves reconstructing 3D poses from 2D keypoint coordinates. Early approaches to monocular 3D pose estimation were based on Hidden Markov Models [27, 67] and graphical models [26, 59]. With the rise of deep learning, Fully Connected Networks [11, 43] were introduced, followed by temporal-based [7, 50] and graph-based [2, 9, 71] convolutional networks, which improved the modeling of sequential dependencies and inter-keypoint relationships.
More recently, transformer-based methods have become the dominant approach, achieving state-of-the-art performance. PoseFormer [85] pioneered this direction, leading to further advancements with models such as [84, 34, 82, 64, 57, 56, 13, 33, 64, 52]. The latest models, KTPFormer [51] integrates kinematics and trajectory priors to improve motion representation, MotionAGFormer [44] introduces a hybrid approach, combining self-attention with graph convolution to better capture both spatial and temporal dependencies, and TCPFormer [37] further enhances temporal modeling by leveraging pose proxies to refine sequential feature learning. Additionally, self-supervised learning has been extensively explored in monocular 3D pose estimation [30, 72, 14, 48, 78, 86].
Although many monocular pose estimation models have been proposed, they are often benchmarked on datasets featuring everyday movements. High-speed sports movements, however, present unique challenges, and the performance of these models on such movements remains unclear. Therefore, in this study, we select the best-performing monocular 2D & 3D models from corresponding benchmark datasets and fine-tune them on the AthletePose3D dataset to evaluate their performance.
@@ 2.3 Validate kinematic from pose estimation
Although models like OpenPose [4], Mediapipe [41], and HRNet [60] are frequently used in sports science and kinematic validation studies, they do not represent the state-of-the-art in 2D pose estimation. Moreover, monocular 3D pose estimation models have received limited attention in validation studies. Therefore, this study evaluates both state-of-the-art monocular 2D and 3D pose estimation models for sports motion kinematic waveform using the newly introduced AthletePose3D dataset.
@@ 3.2 Benchmarking monocular 2D & 3D pose estimation models
To benchmark both monocular 2D and 3D pose estimation models, we considered the best-performing models on representative datasets, taking into account their architectures. These models were fine-tuned and evaluated using the AthletePose3D dataset, which was split into training, validation, and test sets with a 60/20/20 ratio. The training set was used for model training, and the validation set results were reported.